OpenStudy (anonymous):

y'' + (y')^2 + 1 = 0

OpenStudy (anonymous):

are you looking for the double derivative of y?

OpenStudy (anonymous):

I want to know what y is.

OpenStudy (anonymous):

second order diff equations are out of my reach

OpenStudy (anonymous):

Oh, alright, how about integrals, because I got it so far as to have an integral?

OpenStudy (anonymous):

v=y' v'=y'' v'+v^2+1=0 which is saying v'=-(v^2+1) dv/dx=-(v^2+1) dv/(v^2+1)=-dx

OpenStudy (anonymous):

\[\text{ArcTan}[v]\]=-x

OpenStudy (anonymous):

no y?stick issue

OpenStudy (amistre64):

i believe it has something to do with coefficients and "e"

OpenStudy (amistre64):

...but i could be seeing it wrong :)